What is the effective annual yield for an account paying 7% interest compounded monthly? Round your answer to two decimal places and do not include the percent sign.
Answer #1

Given that the rate of interest or annual percentage rate is is 7% (compounded monthly).
When compounding happens monthly, the number of periods in a year will be 12.

Effective annual yield (EAY) is calculated as:
EAY=(1+Rate/Number of periods)^(Number of periods) -1
=(1+7%/12)^(12) -1
=(1.005833333)^(12) -1
=1.072290077 -1
=0.072290077 or 7.23% (Rounded to two decimal places)

Answer: Hence, the effective annual yield is 7.23%